# Causality Guided Representation Learning for Cross-Style Hate Speech Detection

### v1

@@ Introduction
Latent Factor Inference.
@@
$z_{u}=\mu_{u}+\sigma_{u}\odot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$ (3)
@@
$z_{t}^{\text{raw}}=\text{GumbelSoftmax}(\pi_{t}^{\text{raw}},\tau_{t}),$ (4)
@@
Confounder Mitigation
To alleviate the impact of spurious correlation introduced by contextual environment, we design a confounder mitigation mechanism.
Adversarial Confounder Reconstruction.
We employ confounder fusion networks $F$ to control the latent confounder by merging it into the raw variable:
$z_{i}=\sigma_{i}\bigl(F_{i}(z_{i}^{\text{raw}}||z_{u})\bigr),\;i\in\{m,t,s\}$ (5)
@@
$\hat{z}_{u}^{(i)}=D_{i}\bigl(G_{\lambda}(z_{i})\bigr),\;i\in\{m,t,s\}$ (6)
where GRL is a gradient-reversal layer parameterized by the scaling coefficient $\lambda$ so that the feature extractor is updated against the adversary. The resulting adversarial loss is defined by:
@@
Factor-Specific Objective.
@@
$\mathcal{L}_{\text{task}}=\text{CrossEntropy}(\hat{y},y)$ (9)
@@
$o_{ij}=\left(\frac{p_{i}}{\|p_{i}\|_{2}}\cdot\frac{p_{j}}{\|p_{j}\|_{2}}\right)^{2},\;(i,j)\in\{m,t,s,u\}^{2},\;i<j$ (11)
@@
$\mathcal{L}_{\text{orth}}=o_{mt}+o_{ms}+o_{ts}+\lambda_{u}(o_{mu}+o_{tu}+o_{su})$ (12)
@@
$\mathcal{L}_{\text{rec}}=\text{CrossEntropy}(\hat{x},x)$ (16)
Latent Counterfactual Reasoning
@@
$R(z):\;z_{i}^{\text{cf}}=\left\{\begin{array}[]{ll}I-\text{OneHot}(z_{s}),\;i=s\\ \text{Identical}(z_{i}),\;i\in\{m,t,u\}\end{array}\right$. (17)
@@
$\mathcal{L}_{\text{cycle}}=\mathcal{L}_{\text{rec}}(x,\text{Decoder}(R(\text{Encoder}(\text{Decoder}(\hat{h}^{\text{cf}})))))$ (19)
Model Training
@@
$\begin{split}\mathcal{L}=&\lambda_{\text{task}}\mathcal{L}_{\text{task}}+\lambda_{\text{target}}\mathcal{L}_{\text{target}}+\lambda_{\text{style}}\mathcal{L}_{\text{style}}+\lambda_{\text{orth}}\mathcal{L}_{\text{orth}}+\\ &\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}+\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{cf}}\mathcal{L}_{\text{cf}}+\lambda_{\text{cycle}}\mathcal{L}_{\text{cycle}}++\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}\end{split}$ (20)
Where $\mathcal{L}_{\text{KL}}$ terms (one per latent variable) regularize the latent distributions. This streamlined process allows CADET to detect hate speech regardless of its stylistic presentation, providing a principled solution to the challenge of style-invariant detection.
Experiments
@@
Datasets and Evaluation
Our experiments leverage four hate speech datasets: IsHate (Ocampo et al. 2023), Implicit Hate Corpus (IHC), AbuseEval v1.0 (Caselli et al. 2020), and DynaHate as summarized in Table 1. They encompass hate speech in various styles across different platforms, providing a comprehensive evaluation. The detailed information is listed in Appendix A.
We evaluate cross-style generalization under two transfer settings: explicit-to-implicit and implicit-to-explicit hate speech detection. Performance is measured using precision, recall, and macro-F1, averaged over five independent runs to ensure statistical robustness. We benchmark CADET against four categories of baselines: (i) Classic NLP models: BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), and BART (Lewis et al. 2020); (ii) Traditional hate speech detectors: HateBERT (Caselli et al. 2021) and HateXplain (Mathew et al. 2021); (iii) Causal representation methods: PEACE (Sheth et al. 2023) and HateWATCH (Sheth et al. 2024a); and (iv) LLM-based moderation models — ShieldGemma 2B (Zeng et al. 2024) and Llama Prompt Guard 2 (Dubey et al. 2024).
Table 1: Statistics of hate speech datasets.
@@
Table 2: Cross-style Generalization Performance. Bold indicates the best performance and underline the second best.
@@
Cross-Style Generalization
@@
Table 3: Ablation study on each loss component.
@@
Ablation Study and Loss Schedule
@@
We can observe that the full model attains an F1 of 0.99, while all ablated variants show clear degradation, confirming that each objective plays a distinct role in achieving robust, style-invariant detection. Specifically, (i) the counterfactual loss $\mathcal{L}_{\text{cf}}$ has the strongest impact—its removal causes a 24-point F1 drop (0.75), underscoring its importance in enforcing causal consistency across style interventions. Excluding adversarial or reconstruction losses ($\mathcal{L}_{\text{adv}}$, $\mathcal{L}_{\text{rec}}$) yields a 13-point decline, showing that deconfounding and semantic reconstruction are crucial for disentangling hate intent from expression style. Orthogonality and cycle-consistency losses ($\mathcal{L}_{\text{orth}}$, $\mathcal{L}_{\text{cycle}}$) also contribute meaningfully (-7 and -12 points), ensuring that latent factors remain independent and semantically coherent. In contrast, removing $\mathcal{L}_{\text{KL}}$ produces only marginal changes ($\geq$ 1 point), suggesting their roles are largely regularizing. (ii) When multiple objectives are removed, performance drops sharply. Disabling both counterfactual and cycle-consistency terms reduces F1 to 0.78 (-21), indicating their synergy in maintaining semantic stability under style shifts. Removing both adversarial and orthogonality constraints lowers F1 to 0.85 (-14), confirming their complementary roles in deconfounding and latent separation. The most severe degradation (F1 0.70, -30) occurs when all core objectives are excluded, leaving only the basic encoder–decoder structure.
Figure 4: Training loss curve. The training process benefits from the loss-weighting curriculum schedule.
As we mentioned in Section Implementation Details, we adopt a staged training scheme to ensure a competitive training process. We visualize the loss in Figure 4. The total loss initially rises as new objectives—particularly the KL divergence and orthogonality constraints—are introduced, reflecting the model’s adaptation to additional disentanglement pressures. After the reconstruction ramp-up (Epoch 5) and full KL weighting (Epoch 6), the loss steadily decreases, indicating convergence toward a balanced trade-off between reconstruction fidelity and disentanglement. This staged schedule prevents early dominance of complex regularizers, mitigates gradient interference, and ensures smoother, more stable training dynamics for CADET.
@@
Figure 5: Performance on LLM-transformed Hate Speech. CADET is robust under controlled style transformations.
Figure 6: t-SNE visualization for latent factors. CADET isolates hate motivation from platform-dependent target and style.
Controlled Style Transformations
RQ3: Can CADET maintain high detection performance when hateful intent is rephrased from explicit to implicit style? To examine this, we constructed a controlled evaluation set by systematically transforming explicit hate speech into implicit equivalents using GPT-4. A carefully designed prompt (Appendix D) ensured that the hateful intent and target group remained unchanged while only the stylistic surface, such as the presence of slurs, direct insults, or overt hostility, was altered. From 1,000 explicit hate tweets drawn across datasets, we retained 375 high-quality implicit rephrasings after human verification for semantic fidelity. Each model was trained solely on explicit hate speech and evaluated on these implicit counterparts, mimicking a realistic deployment scenario where detectors face reworded, subtler expressions of hate.
@@
Latent Factor Visualization
@@
Figure 7: Case Study.
Conclusion
@@ Appendix A Dataset Details
We consider four different dataset in our work: (i) IsHate (Ocampo et al. 2023) comprises approximately 29,116 social media posts specifically curated to study subtle and implicit hate. This dataset is particularly valuable for our causal analysis as it contains substantial numbers of both explicit and implicit hate examples, each annotated with fine-grained labels indicating explicitness and subtlety. These annotations directly correspond to the Content Strategy (S) variable in our causal graph. (ii) Implicit Hate Corpus (IHC) (ElSherief et al. 2021) offers 22,584 tweets from extremist groups, with 6,346 labeled as implicit hate. IHC’s granular annotations not only distinguish between implicit and explicit hate but also provide implied statements that articulate the underlying hateful meaning—making it an ideal testbed for evaluating CADET’s ability to extract invariant hateful intent. (iii) AbuseEval v1.0 (Caselli et al. 2020) enhances the OLID/OffensEval Twitter dataset with explicit vs. implicit labeling for both offensive and abusive content. Its two-dimensional annotation scheme allows us to isolate examples by explicitness for our cross-style experiments. (iv) DynaHate (Vidgen et al. 2021) features challenging hate speech examples created through a human-and-model-in-the-loop process, resulting in a dataset enriched with more covert hateful expressions designed to evade detection. For our cross-style experiments, we heuristically separate DynaHate posts into explicit and implicit categories based on the presence of overt slurs and epithets.
@@ Appendix B Baseline Details
  • BERT (Devlin et al. 2019), RoBERTa (Liu et al. 2019), and BART (Lewis et al. 2020) serve as strong classic NLP baselines. These Transformer-based encoders and encoder–decoder models capture contextual semantics effectively but lack explicit mechanisms for causal or style-invariant reasoning. They represent the standard supervised fine-tuning paradigm widely adopted in text classification.
  • HateBERT (Caselli et al. 2021) adapts BERT via domain-specific pre-training on banned Reddit communities containing toxic discourse. This enhances sensitivity to abusive language but provides limited robustness to stylistic or contextual shifts.
  • HateXplain (Mathew et al. 2021) extends BERT with an explainability module and multi-task learning using human rationales. While it improves interpretability and fairness, it does not explicitly disentangle style from underlying hate intent.
  • PEACE (Sheth et al. 2023) introduces causality-inspired cues—sentiment and aggression—as auxiliary features for hate detection. Although it partially captures invariant representations, it lacks explicit counterfactual modeling of stylistic variations.
  • HateWATCH (Sheth et al. 2024a) employs weakly supervised causal disentanglement through contrastive learning to separate invariant hateful content from group- or domain-specific signals. It is conceptually closest to our framework but does not perform structured style interventions or counterfactual reasoning.
  • ShieldGemma 2B (Zeng et al. 2024) and Llama Prompt Guard 2 (Dubey et al. 2024) represent recent LLM-based moderation models, considering the prevailing role of LLMs in various domains (Li et al. 2025; Zhao et al. 2025a, b). These systems leverage instruction-tuned large language models for safety classification, trained on extensive curated moderation datasets. While highly capable in open-ended text understanding, they are primarily optimized for general safety filtering rather than fine-grained cross-style hate speech reasoning. As such, they offer a strong but stylistically uncalibrated comparison point for evaluating causal generalization.
@@ Appendix D Experiment Details
Using this prompt, we transformed 1,000 explicit hate tweets sampled from the datasets described in Section Controlled Style Transformations. After manual inspection for semantic fidelity, we retained 375 high-quality implicit rephrasings in which the hateful meaning and target identity were faithfully preserved.
For evaluation, each model was trained exclusively on explicit hate speech and subsequently tested on these implicit rephrasings, simulating a realistic deployment scenario where a detector trained on explicit data encounters implicit manifestations of the same underlying hate. To further validate the integrity of the transformation process, we also generated a parallel set of examples in which hateful statements were rephrased into non-hateful counterparts. This additional set was used solely for quality assurance—to confirm that models were not simply flagging any rephrased text as hateful—and was excluded from quantitative evaluation.

### v2

@@ 4.2.1. Text Encoding
4.2.2. Latent Factor Inference.
@@
(3) $z_{u}=\mu_{u}+\sigma_{u}\odot\epsilon,\quad\epsilon\sim\mathcal{N}(0,I),$
@@
(4) $z_{t}^{\text{raw}}=\text{GumbelSoftmax}(\pi_{t}^{\text{raw}},\tau_{t}),$
@@
4.3. Confounder Mitigation
To alleviate the impact of spurious correlations introduced by the context, we design a confounder-mitigation mechanism.
4.3.1. Adversarial Confounder Reconstruction.
We employ confounder fusion networks $F$ to control the latent confounder by merging it into the raw variables:
(5) $z_{i}=\sigma_{i}\bigl(F_{i}(z_{i}^{\text{raw}}||z_{u})\bigr),\;i\in\{m,t,s\}$
@@ 4.3. Confounder Mitigation
(6) $\hat{z}_{u}^{(i)}=D_{i}\bigl(G_{\lambda}(z_{i})\bigr),\;i\in\{m,t,s\}$
where $G_{\lambda}$ is a gradient-reversal layer parameterized by the scaling coefficient $\lambda$ so that the feature extractor is updated against the adversary. The resulting adversarial loss is defined by:
@@
4.3.2. Factor-Specific Objective.
@@
(9) $\mathcal{L}_{\text{task}}=\text{CrossEntropy}(\hat{y},y)$
@@ 4.3.3. Orthogonality Constraints
(11) $o_{ij}=\left(\frac{p_{i}}{\|p_{i}\|_{2}}\cdot\frac{p_{j}}{\|p_{j}\|_{2}}\right)^{2},\;(i,j)\in\{m,t,s,u\}^{2},\;i<j$
@@
(12) $\mathcal{L}_{\text{orth}}=o_{mt}+o_{ms}+o_{ts}+\lambda_{u}(o_{mu}+o_{tu}+o_{su})$
@@ 4.3.4. Input Reconstruction
(16) $\mathcal{L}_{\text{rec}}=\text{CrossEntropy}(\hat{x},x)$
4.4. Latent Counterfactual Reasoning
@@ 4.4.1. Style Intervention
(17) $R(z):\;z_{i}^{\text{cf}}=\left\{\begin{array}[]{ll}I-\text{OneHot}(z_{s}),\;i=s\\ \text{Identical}(z_{i}),\;i\in\{m,t,u\}\end{array}\right$.
@@ 4.4.3. Cycle Consistency
(19) $\mathcal{L}_{\text{cycle}}=\mathcal{L}_{\text{rec}}(x,\text{Decoder}(R(\text{Encoder}(\text{Decoder}(\hat{h}^{\text{cf}})))))$
4.5. Model Training
@@
(20) $\begin{split}\mathcal{L}=\ &\lambda_{\text{task}}\mathcal{L}_{\text{task}}+\lambda_{\text{target}}\mathcal{L}_{\text{target}}+\lambda_{\text{style}}\mathcal{L}_{\text{style}}+\lambda_{\text{orth}}\mathcal{L}_{\text{orth}}+\\ &\lambda_{\text{adv}}\mathcal{L}_{\text{adv}}+\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{cf}}\mathcal{L}_{\text{cf}}+\lambda_{\text{cycle}}\mathcal{L}_{\text{cycle}}++\lambda_{\text{KL}}\mathcal{L}_{\text{KL}}\end{split}$
where $\mathcal{L}_{\text{KL}}$ terms regularize the latent distributions.
This streamlined process allows CADET to detect hate speech regardless of its stylistic presentation, providing a principled solution to the challenge of style-invariant detection.
5. Experiments
@@
5.1. Datasets and Evaluation
Our experiments leverage four hate speech datasets: IsHate (Ocampo et al., 2023b), Implicit Hate Corpus (IHC), AbuseEval v1.0 (Caselli et al., 2020), and DynaHate (Vidgen et al., 2021) as summarized in Table 1. They encompass hate speech in various styles across different platforms, providing a comprehensive evaluation. The detailed information is listed in Appendix A.
We evaluate cross-style hate speech detection across two settings: explicit-to-implicit and implicit-to-explicit. Performance is measured using precision, recall, and macro-F1, averaged over five independent runs to ensure statistical robustness. We benchmark CADET against four categories of baselines: (i) Classic NLP models: BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), and DistilBERT (Sanh, 2019); (ii) Traditional hate speech detectors: HateBERT (Caselli et al., 2021) and HateXplain (Mathew et al., 2021); (iii) Causal representation methods: PEACE (Sheth et al., 2023) and HateWATCH (Sheth et al., 2024a); and (iv) LLM-based moderation models – ShieldGemma 2B (Zeng et al., 2024) and Llama Prompt Guard 2 (Dubey et al., 2024).
Table 1. Statistics of hate speech datasets.
@@ 5.1. Datasets and Evaluation
Table 2. Cross-style Generalization Performance. Bold indicates the best performance and underline the second best.
@@ 5.2. Implementation Details
5.3. Cross-Style Generalization
@@
Table 3. Ablation study on each loss component.
@@
5.4. Ablation Study and Loss Schedule
@@
We can observe that the full model attains an F1 of 0.99, while all ablated variants show clear degradation, confirming that each objective plays a distinct role in achieving robust, style-invariant detection. Specifically, (i) the counterfactual loss $\mathcal{L}_{\text{cf}}$ has the strongest impact–its removal causes a 24-point F1 drop (0.75), underscoring its importance in enforcing causal consistency across style interventions. Excluding adversarial or reconstruction losses ($\mathcal{L}_{\text{adv}}$, $\mathcal{L}_{\text{rec}}$) yields a 13-point decline, showing that deconfounding and semantic reconstruction are crucial for disentangling hate intent from expression style. Orthogonality and cycle-consistency losses ($\mathcal{L}_{\text{orth}}$, $\mathcal{L}_{\text{cycle}}$) also contribute meaningfully (-7 and -12 points), ensuring that latent factors remain independent and semantically coherent. In contrast, removing $\mathcal{L}_{\text{KL}}$ produces only marginal changes ($\geq$ 1 point), suggesting their roles are largely regularizing. (ii) When multiple objectives are removed, performance drops sharply. Disabling both counterfactual and cycle-consistency terms reduces F1 to 0.78 (-21), indicating their synergy in maintaining semantic stability under style shifts. Removing both adversarial and orthogonality constraints lowers F1 to 0.85 (-14), confirming their complementary roles in deconfounding and latent separation. The most severe degradation (F1 0.70, -30) occurs when all core objectives are excluded, leaving only the basic encoder–decoder structure.
As we mentioned in Section 5.2, we adopt a staged training scheme to ensure a competitive training process. We visualize the loss in Figure 4. The total loss initially rises as new objectives, particularly the KL divergence and orthogonality constraints, are introduced, reflecting the model’s adaptation to additional disentanglement pressures. After the reconstruction ramp-up (Epoch 5) and full KL weighting (Epoch 6), the loss steadily decreases, indicating convergence toward a balanced trade-off between reconstruction fidelity and disentanglement. This staged schedule prevents early dominance of complex regularizers, mitigates gradient interference, and ensures smoother, more stable training dynamics for CADET.
Figure 4. Training loss curve. The training process benefits from the loss-weighting curriculum schedule.
@@
5.5. Controlled Style Transformations
RQ3: Can CADET maintain high detection performance when hateful intent is rephrased from explicit to implicit style? To examine this, we constructed a controlled evaluation set by systematically transforming explicit hate speech into implicit equivalents using GPT-4. A carefully designed prompt (Appendix D.1) ensured that the hateful intent and target group remained unchanged while only the stylistic surface, such as the presence of slurs, direct insults, or overt hostility, was altered. From 1,000 explicit hate tweets drawn across datasets, we retained 375 high-quality implicit rephrasings after human verification for semantic fidelity. Each model was trained solely on explicit hate speech and evaluated on these implicit counterparts, mimicking a realistic deployment scenario where detectors face reworded, subtler expressions of hate.
Figure 5. Performance on LLM-transformed Hate Speech. CADET is robust under controlled style transformations.
Figure 6. t-SNE visualization for latent factors. CADET isolates hate motivation from platform-dependent target and style.
@@ 5.5. Controlled Style Transformations
5.6. Latent Factor Visualization
@@ 5.7. Case Study
Figure 7. Case Study.
6. Conclusion
@@
Acknowledgments
This work is supported by the National Science Foundation (NSF) under grants #2311716, #2435886 and SaTC #2335666.
@@ Appendix A Dataset Details
We compile an Implicit-Explicit Hate Speech Corpus (IE-HSC) [^2] Dataset: https://huggingface.co/datasets/Shuwan/cadet-datasets via four different datasets in our work: (i) IsHate (Ocampo et al., 2023b) comprises approximately 29,116 social media posts specifically curated to study subtle and implicit hate. This dataset is particularly valuable for our causal analysis as it contains substantial numbers of both explicit and implicit hate examples, each annotated with fine-grained labels indicating explicitness and subtlety. These annotations directly correspond to the Content Strategy (S) variable in our causal graph. (ii) Implicit Hate Corpus (IHC) (ElSherief et al., 2021) offers 22,584 tweets from extremist groups, with 6,346 labeled as implicit hate. IHC’s granular annotations not only distinguish between implicit and explicit hate but also provide implied statements that articulate the underlying hateful meaning—making it an ideal testbed for evaluating CADET’s ability to extract invariant hateful intent. (iii) AbuseEval v1.0 (Caselli et al., 2020) enhances the OLID/OffensEval Twitter dataset with explicit vs. implicit labeling for both offensive and abusive content. Its two-dimensional annotation scheme allows us to isolate examples by explicitness for our cross-style experiments. (iv) DynaHate (Vidgen et al., 2021) features challenging hate speech examples created through a human-and-model-in-the-loop process, resulting in a dataset enriched with more covert hateful expressions designed to evade detection. For our cross-style experiments, we heuristically separate DynaHate posts into explicit and implicit categories based on the presence of overt slurs and epithets.
@@ Appendix B Baseline Details
  • BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), and DistilBERT (Sanh, 2019) serve as strong classic NLP baselines. These Transformer-based encoders and encoder–decoder models capture contextual semantics effectively but lack explicit mechanisms for causal or style-invariant reasoning. They represent the standard supervised fine-tuning paradigm widely adopted in text classification.
  • HateBERT (Caselli et al., 2021) adapts BERT via domain-specific pre-training on banned Reddit communities containing toxic discourse. This enhances sensitivity to abusive language but provides limited robustness to stylistic or contextual shifts.
  • HateXplain (Mathew et al., 2021) extends BERT with an explainability module and multi-task learning using human rationales. While it improves interpretability and fairness, it does not explicitly disentangle style from underlying hate intent.
  • PEACE (Sheth et al., 2023) introduces causality-inspired cues–sentiment and aggression–as auxiliary features for hate detection. Although it partially captures invariant representations, it lacks explicit counterfactual modeling of stylistic variations.
  • HateWATCH (Sheth et al., 2024a) employs weakly supervised causal disentanglement through contrastive learning to separate invariant hateful content from group- or domain-specific signals. It is conceptually closest to our framework but does not perform structured style interventions or counterfactual reasoning.
  • ShieldGemma 2B (Zeng et al., 2024) and Llama Prompt Guard 2 (Dubey et al., 2024) represent recent LLM-based moderation models, considering the prevailing role of LLMs in various domains (Li et al., 2025; Zhao et al., 2025a, b). These systems leverage instruction-tuned large language models for safety classification, trained on extensively curated moderation datasets. While highly capable in open-ended text understanding, they are primarily optimized for general safety filtering rather than fine-grained cross-style hate speech reasoning. As such, they offer a strong but stylistically uncalibrated comparison point for evaluating causal generalization.
@@ D.1. Illustration of Prompt
Using this prompt, we transformed 1,000 explicit hate tweets sampled from the datasets described in Section LABEL:sec:datasets. After manual inspection for semantic fidelity, we retained 375 high-quality implicit rephrasings in which the hateful meaning and target identity were faithfully preserved.
For evaluation, each model was trained exclusively on explicit hate speech and subsequently tested on these implicit rephrasings, simulating a realistic deployment scenario where a detector trained on explicit data encounters implicit manifestations of the same underlying hate. To further validate the integrity of the transformation process, we also generated a parallel set of examples in which hateful statements were rephrased into non-hateful counterparts. This additional set was used solely for quality assurance–to confirm that models were not simply flagging any rephrased text as hateful–and was excluded from quantitative evaluation.